OpenStudy (anonymous):

a communications satellite transmits a radio wave at a frequency of 5.0x10^9 Hz. What is the signal's wavelength? Assume the wave travels in a vacuum. show your work.

OpenStudy (astrophysics):

Well we know that the velocity with be the speed of light because we're dealing with EMR so we can use the formula \[c = f \times \lambda \implies \lambda = \frac{ c }{ f }\]

OpenStudy (shamim):

Given c=3*10^8m/s f=5*10^9Hz Lamda=?